# An interpretation of the Brownian bridge as a physics-informed prior for the Poisson equation

### v1

@@ Abstract
Physics-informed machine learning is one of the most commonly used methods for fusing physical knowledge in the form of partial differential equations with experimental data. The idea is to construct a loss function where the physical laws take the place of a regularizer and minimize it to reconstruct the underlying physical fields and any missing parameters. However, there is a noticeable lack of a direct connection between physics-informed loss functions and an overarching Bayesian framework. In this work, we demonstrate that Brownian bridge Gaussian processes can be viewed as a softly-enforced physics-constrained prior for the Poisson equation. We first show equivalence between the variational form of the physics-informed loss function for the Poisson equation and a kernel ridge regression objective. Then, through the connection between Gaussian process regression and kernel methods, we identify a Gaussian process for which the posterior mean function and physics-informed loss function minimizer agree. This connection allows us to probe different theoretical questions, such as convergence and behavior of inverse problems. We also connect the method to the important problem of identifying model-form error in applications.
Keywords: Scientific machine learning, inverse problems, Poisson equation, Gaussian process regression, reproducing kernel Hilbert spaces
@@ 1 Introduction
$\left\{\begin{split}\Delta u+q&=0\quad\mathrm{on}\quad\Omega\subset\mathbb{R}^{d}\\ u&=0\quad\mathrm{on}\quad\partial\Omega.\end{split}\right.$ (1.1)
We assume that the source term $q$ is sufficiently regular and $\Omega=[0,1]^{d}$ so that eq. (1.1) is well-posed in the classical sense, and the solution, $u^{0}$, lives in the usual space $H_{p}\coloneqq\{u\in H^{1}(\Omega)\cap H^{2}(\Omega):u=0\>\mathrm{on}\>\partial\Omega\}$. Here by $H^{1}(\Omega)$ and $H^{2}(\Omega)$, we are referring to Sobolev spaces. That is, given $\tau\in\mathbb{N}$ denote the Sobolev space of square integrable functions on $\Omega$ with square-integrable weak derivatives up to order $\tau$ by $H^{\tau}(\Omega)$:
@@
for a multi-index $\alpha$.
Further, we assume we have access to some measurement data appearing in the typical way $y_{i}=R_{i}u+\gamma_{i}$, $i=1,\dots,n$, where $y_{i}\in\mathbb{R}$ are the individual measurements, and $\gamma_{i}$ represents zero-mean additive noise to the measurement. Each functional $R_{i}:H\to\mathbb{R}$ is called a measurement operator and describes the process in which data are generated. At the moment, we will assume that $R_{i}$ is continuous and linear, and we work under the assumption that the measurement noise $\gamma_{i}$ follows a zero-mean i.i.d. Gaussian model. Nonlinear measurements and non-Gaussian noise are more involved in the setting of Gaussian process (GP) regression, but can be incorporated. Here, we are interested in the derivation of a Bayesian approach for solving the Poisson equation, which treats the PDE as prior information. The application we have in mind is the inverse problem, where $q$ is unknown and needs to be identified.
By now, the idea of physics-informed machine learning [30] has become a standard framework for solving this kind of problem. In the simplest case with noise-free point observations, this is approached as an optimization problem. Specifically, we approximate $u$ with a class of functions $\hat{u}(\cdot;\theta)$ with trainable parameters $\theta$. If $\hat{u}$ is a neural network, then we are in the regime of well-known physics-informed neural networks (PINNs) [39]. However, the choice of parameterization is not restricted to a PINN, e.g., $\hat{u}$ could be a truncated basis expansion.
Under this framework, we derive a physics-informed loss function from the PDE for training. Rather than starting with a parameterization, we will derive the loss function in the function space setting. The first step is to cast solving eq. (1.1) as an optimization problem. As we are working with the Poisson equation, we have access to a variational formulation through means of Dirichelt’s principle [12]:
$\left\{\begin{split}\min_{u\in H_{p}}\quad&E(u)\coloneqq\int_{\Omega}\frac{1}{2}\|\nabla u\|^{2}-qu\>d\Omega\\ \textrm{s.t.}\quad&u=0\quad\mathrm{on}\quad\partial\Omega.\end{split}\right.$ (1.2)
We will refer to $E(u)$ in the above as the energy functional. Although it is more common to use the integrated square residual of eq. (1.1) in physics-informed machine learning literature, this form of the energy functional is in some sense better behaved when compared to the integrated square residual, due to the fact that it yields a convex optimization problem. Variational forms are sometimes used when training PINNs, e.g., [31, 6].
@@
Equations (1.2) and eq. (1.3) are combined to construct the training loss function:
@@
where $\eta>0$ is a regularization parameter chosen to balance contributions from the data and physics. A loss balancing approach may be used to adjust the learning rate as the PINN trains, although in some cases we may hand-pick $\eta$ [46]. The field reconstruction problem is then solved by minimizing this loss function.
In the more difficult case with measurement noise, the field reconstruction problem can easily become ill-posed, so a Bayesian approach is desirable. For PINNs, perhaps the most direct method for this is Bayesian-PINNs [48]. In Bayesian-PINNs, a standard normal prior is placed on the network parameters, and the square residuals of the physics are used to construct an additional likelihood term by introducing fictitious noise under a Gaussian measurement model. Similarly to Bayesian-PINNs, [36] rethinks how PINNs are trained with noisy data. The PDE solution is treated as a GP with prior mean function and covariance parameterized by neural networks. This essentially adds a kernel-structure to the PINN weights. However, there is a notable lack of a direct connection between the physics-informed loss function and an overarching Bayesian interpretation. The goal of this work is to establish a connection between this physics-informed loss function and a GP regression method.
@@
There are other works which go about incorporating physics into the Bayesian field reconstruction problem through different means, usually under a GP framework. If the physics are linear, as in our case, it is possible to define GP priors from the physics by careful consideration of the covariance kernel. For example, this idea can be found in [40, 7] where the covariance kernel is constructed using a numerical solver for the PDE. There is also [26], which is restricted only to linear ODEs with constant coefficients. Nonlinear PDEs can be handled by promoting the physics to the likelihood [15] and using a standard GP prior, e.g., the square exponential kernel. The maximum a posteriori (MAP) estimate is then taken as the solution to the PDE. A similar work can be found in [16]. This work again treats the PDE solution as a GP prior. The solution is identified by minimizing the reproducing kernel Hilbert space (RKHS) norm of the prior covariance constrained on the PDE residuals on a predefined grid. This of course adds additional assumptions through a regularizer which enforces smoothness and may not directly represent the underlying physics. Also, in both methods only a deterministic answer is given. One may also argue that treating the physics as a likelihood is philosophically unsatisfying under the Bayesian treatment.
Of note is the work in [2]. Through application of Mercer’s theorem to construct a particular covariance kernel, GPs are defined whose samples are exact solutions to linear PDEs. This idea also inspired the physics-consistent neural networks as an alternative to PINNs [42]. However, this also introduces an interesting fundamental mismatch between the GP regression solution and restrictions imposed by the PDE: GP samples are a.s. not in the RKHS of the covariance, see Theorem 2.2. In particular, the posterior mean function will reside in the RKHS of the prior covariance, for which the GP samples cannot live in. By necessity the GP posterior mean function will a.s. not be a solution to the PDE. In this work, we construct our GP prior informed from the Poisson equation which behaves much differently. Namely, we derive a GP with the opposite behavior.
@@ 1.1 Contributions
  1. (i)
    We show the energy-based physics-informed loss function for the Poisson equation yields the MAP estimator of a GP regression scheme with the Brownian bridge as the prior.
  2. (ii)
    We derive a finite-dimensional representation of the prior for use in applications. The representation places the mesh on $L^{2}(\Omega)$ rather than $\Omega$ as is typical in GP regression.
  3. (iii)
    We prove this MAP estimator, and hence the function which solves eq. (1.4), converges to the ground truth in the large-data limit. Convergence holds even in the presence of significant model-form error.
  4. (iv)
    By tuning an additional hyperparameter of the prior, we connect the method to the problem of identifying model-form error. We show this hyperparameter, which controls the prior variance, is sensitive to model-form error by enforcing the physics as a soft constraint. The hyperparameter also causes the variance of the approximation to $q$ to adjust in the context of inverse problems.
@@ 1.2 Outline
The paper is organized as follows. In Section 2, we provide the necessary background on Gaussian process regression and kernel ridge regression. Section 3 establishes the connection between the physics-informed loss function of eq. (1.4) and the Brownian bridge GP. We do so by showing the loss function is the related kernel method objective, from which we deduce it is the MAP estimate of the corresponding GP regression. In 1D, we also prove the result in the setting of infinite-dimensional Bayesian inverse problems. That is, we show eq. (1.4) is the MAP estimate of the posterior obtained when starting with the Brownian bridge as a Gaussian measure on $L^{2}([0,1])$. Some analysis of the method is explored in Section 4. Here, we state the regularity of the prior and establish convergence conditions for the MAP estimate. We also derive a finite-dimensional approximation to the prior. Finally, in Section 5, we connect the method to the problem of model-form error identification. We demonstrate that the posterior of the inverse problem adjusts according to error in the specified physical model.
@@ 2 Preliminaries
We provide some necessary background information on GPs and RKHSs required in this work. In Appendix A, we also provide a background on the theory of Gaussian measures, which, while used somewhat, is not the main focus in this work.
Definition 2.1 (Reproducing kernel Hilbert space).
Let $k$ be a positive definite kernel on $\Omega\times\Omega$. A Hilbert space $H_{k}$ on $\Omega$ equipped with inner product $\langle\cdot,\cdot\rangle_{H_{k}}$ is said to be a reproducing kernel Hilbert space if the following two properties hold:
  1. (i)
@@
Property (ii) of Definition 2.1 is called the reproducing property, and the kernel defining the RKHS is called the reproducing kernel. The RKHS is uniquely determined by the positive-definite kernel that defines it, and the reverse is also true. This results from the Moore-Aronszajn theorem [4], which states that every positive definite kernel $k$ is associated with a unique RKHS $H_{k}$ for which $k$ is the reproducing kernel. Likewise, for every RKHS $H_{k}$ by definition there exists a unique positive definite kernel which satisfies properties (i) and (ii) of Definition 2.1. In this way, there is a one-to-one correspondence. One can show that given a positive definite kernel $k$ and its RKHS, each $f\in H_{k}$ can be written as $f=\sum_{i=1}^{\infty}\alpha_{i}k(\cdot,x_{i})$ for some $(\alpha_{i})_{i=1}^{\infty}\subset\mathbb{R}$, $(x_{i})_{i=1}^{\infty}\subset\Omega$ and $\|f\|_{H_{k}}<\infty$, where $\|f\|_{H_{k}}^{2}\coloneqq\sum_{i,j=1}^{\infty}c_{i}c_{j}k(x_{i},x_{j})$. It is therefore easy to verify that the functions in the RKHS inherit the properties of $k$, e.g., smoothness.
@@
[…]
The assumptions on $k$ imply that $L_{k}$ is a self-adjoint, positive operator, and thus has spectral decomposition
$(L_{k}u)(x)=\sum_{n\in\mathbb{N}}\lambda_{n}\langle u,\psi_{n}\rangle\psi_{n}(x),$
where $(\lambda_{n},\psi_{n})_{n=1}^{\infty}$ is the eigensystem of $L_{k}$, i.e.
$L_{k}\psi_{n}=\lambda_{n}\psi_{n},$ (2.2)
@@
Theorem 2.1 (Mercer’s Theorem [43]).
Let $k:\Omega\times\Omega\to\mathbb{R}$ be a continuous, positive-definite kernel, and $L_{k}$ and $(\lambda_{n},\psi_{n})_{n=1}^{\infty}$ be as given in eq. (2.1) and eq. (2.2), respectively. Then,
@@
$H_{k}=\left\{u\in L^{2}(\Omega):\sum_{n\in\mathbb{N}}\frac{1}{\lambda_{n}}\langle u,\psi_{n}\rangle<\infty\right\},$
@@
for $u,v\in H_{k}$. Hence the RKHS-norm can be expressed as $\|u\|_{H_{k}}^{2}=\sum_{n=1}^{\infty}\lambda_{n}^{-1}\langle u,\psi_{n}\rangle^{2}$. This representation is useful to us later when constructing the RKHS associated with the physics-informed loss function.
@@
Definition 2.2 (Gaussian process).
@@
Theorem 2.2 (Theorem 3.1 [29]).
Assume we have data given by (2.3) and a GP prior $u\sim\mathcal{GP}(m,k)$. Then the posterior follows $u|y\sim\mathcal{GP}(\tilde{m},\tilde{k})$, where
$\tilde{m}(x)$ $\coloneqq m(x)+k_{xX}(K_{XX}+\sigma^{2}I_{n})^{-1}(y-m_{X}),\quad x\in\Omega$ (2.4)
$\tilde{k}(x,x^{\prime})$ $\coloneqq k(x,x^{\prime})-k_{xX}(K_{XX}+\sigma^{2}I_{n})^{-1}k_{Xx^{\prime}},\quad x,x^{\prime}\in\Omega,$ (2.5)
@@
Kernel ridge regression (KRR), or regularized least squares [13], is closely related to GP regression. Given data in eq. (2.3), the objective of KRR is to solve the following interpolation problem
$u^{*}=\operatorname*{arg\,min}_{u\in H_{k}}\frac{1}{n}\sum_{i=1}^{n}\left(u(x_{i})-y_{i}\right)^{2}+\eta\|u\|_{H_{k}}^{2},$ (2.6)
where $\eta\geq 0$ is the regularization parameter. The inclusion of the RKHS norm in the objective function serves as a regularizer which enforces the class of functions which fit the data, while simultaneously smoothing the fit. It is known that $u$ becomes smoother as $\|u\|_{H_{k}}$ gets smaller. Specifying the kernel which defines the KRR objective eq. (2.6) effectively enforces a prior on the fit. As with the GP regression posterior mean function, the solution to eq. (2.6) is also unique:
Theorem 2.3 (Theorem 3.4 [29]).
Let $\eta>0$. Then the unique solution to eq. (2.6) is
$u^{*}(x)=k_{xX}(K_{XX}+n\eta I_{n})^{-1}y=\sum_{i=1}^{n}\alpha_{i}k(x,x_{i}),\quad x\in\Omega,$
where $k_{xX}=k_{Xx}^{T}\coloneqq(k(x,x_{1}),\dots,k(x,x_{n}))^{T}$ and $(\alpha_{1},\dots,\alpha_{n})^{T}=(K_{XX}+n\eta I_{n})^{-1}d$. Further, if the matrix $K_{XX}$ is invertible, then the coefficients $\alpha_{i}$ are unique.
In [29], the relationship between GP regression and KRR is discussed in great detail. In a certain sense, GP regression can be viewed as the Bayesian interpretation of KRR. Notably, under mild conditions, the KRR solution and GP posterior mean function are equivalent.
@@
Let $k:\Omega\times\Omega\to\mathbb{R}$ be a positive definite kernel, and eq. (2.3) be training data. If $\sigma^{2}=n\lambda$, then $\tilde{m}=u^{*}$, where $\tilde{m}$ is the GP posterior mean function and $u^{*}$ is the unique KRR solution, given by eq. (2.4) and eq. (2.6), respectively.
The equivalence between the GP posterior mean and KRR solution helps to establish much of the behavior involved with GP regression in terms of the RKHS of the prior covariance kernel. For example, it is immediate from Proposition 2.1 that the GP posterior mean function lives in the RKHS of the prior, meaning that the behavior of the posterior mean is inherited from the specified prior covariance. The last important property we need is the fact that GP sample paths a.s. do not belong to the prior RKHS, which is a consequence of Driscol’s zero-one law [18].
Proposition 2.2 (Corollary 4.10 [29]).
Let $k:\Omega\times\Omega\to\mathbb{R}$ be a positive definite kernel and $H_{k}$ be the corresponding RKHS. Let $u\sim\mathcal{GP}(m,k)$ where $m\in H_{k}$. If $H_{k}$ is infinite-dimensional, then $u\in H_{k}$ with probability 00.

### v2

@@ Abstract
Many inverse problems require reconstructing physical fields from limited and noisy data while incorporating known governing equations. A growing body of work within probabilistic numerics formalizes such tasks via Bayesian inference in function spaces by assigning a physically meaningful prior to the latent field. In this work, we demonstrate that Brownian bridge Gaussian processes can be viewed as a softly-enforced physics-constrained prior for the Poisson equation. We first show equivalence between the variational problem associated with the Poisson equation and a kernel ridge regression objective. Then, through the connection between Gaussian process regression and kernel methods, we identify a Gaussian process for which the posterior mean function and the minimizer to the variational problem agree, thereby placing this PDE-based regularization within a fully Bayesian framework. This connection allows us to probe different theoretical questions, such as convergence and behavior of inverse problems. We then develop a finite-dimensional representation in function space and prove convergence of the projected prior and resulting posterior in Wasserstein distance. Finally, we connect the method to the important problem of identifying model-form error in applications, providing a diagnostic for model misspecification.
Keywords: Probabilistic numerics, scientific machine learning, inverse problems, Poisson equation, Gaussian process regression, reproducing kernel Hilbert spaces
@@ 1 Introduction
$\left\{\begin{split}\Delta u+q&=0\qquad\mathrm{on}\qquad\Omega\subset\mathbb{R}^{d}\\ u&=0\qquad\mathrm{on}\qquad\partial\Omega.\end{split}\right$. (1.1)
We assume that the source term $q$ is sufficiently regular and $\Omega=[0,1]^{d}$ so that eq. (1.1) possesses a weak solution $u^{0}\in H^{1}_{0}(\Omega)$, which denotes the subset of $H^{1}(\Omega)$ of functions which are also zero on the boundary. Here by $H^{1}(\Omega)$, we are referring to the first-order Sobolev space. That is, given $\tau\in\mathbb{N}$ denote the Sobolev space of square integrable functions on $\Omega$ with square-integrable weak derivatives up to order $\tau$ by $H^{\tau}(\Omega)$:
@@
for a multi-index $\alpha$. The exact nature of $q$ depends on the dimension, and is discussed later.
Further, we assume we have access to some measurement data appearing in the typical way $y_{i}=R_{i}u+\gamma_{i}$, $i=1,\dots,n$, where $y_{i}\in\mathbb{R}$ are the individual measurements, and $\gamma_{i}$ represents zero-mean additive noise to the measurement. Each functional $R_{i}:H\to\mathbb{R}$ is called a measurement operator and describes the process that generates the data. At the moment, we assume that $R_{i}$ is continuous and linear and that the measurement noise $\gamma_{i}$ follows a zero-mean i.i.d. Gaussian model. Nonlinear measurements and non-Gaussian noise are more involved in the setting of Gaussian process (GP) regression, but can be incorporated. Here, we are interested in the derivation of a Bayesian approach for solving the Poisson equation, which treats the PDE as prior information. The application we have in mind is the inverse problem, where $q$ is unknown and needs to be identified.
The Poisson model underlies a wide range of applications including electrostatics and gravitation, steady-state heat conduction and diffusion, potential theory and pressure projection in incompressible flow, and imaging tasks such as Poisson image editing, shape-from-shading, and variational denoising. In many of these settings the quantity of interest (the source, a coefficient, or boundary data) is only indirectly observed through linear functionals of the state, so the measurement model above is natural. Casting the PDE itself as prior information places our method within probabilistic numerics [hennig2015probabilistic]: the elliptic operator furnishes a Gaussian prior, via its Green’s function, that encodes regularity and boundary conditions, and Bayesian updating then yields both estimators and calibrated uncertainty. The result is a probability distribution over the solution of the PDE.
We treat solving the PDE as an inverse problem by identifying a suitable loss function. The first step is to cast solving eq. (1.1) as an optimization problem. As we are working with the Poisson equation, we have access to a variational formulation through means of Dirichelt’s principle [brezis2011functional]:
$\left\{\begin{split}\min_{u\in H^{1}_{0}(\Omega)}\qquad&E(u)\coloneqq\int_{\Omega}\frac{1}{2}\|\nabla u\|^{2}-qu\>d\Omega\\ \textrm{s.t.}\qquad&u=0\qquad\mathrm{on}\qquad\partial\Omega.\end{split}\right$. (1.2)
We will refer to $E(u)$ in the above as the energy functional. Although it is more common to use the integrated square residual of eq. (1.1) in much of the related literature, e.g., in physics-informed neural networks (PINNs) [raissi2019physics], this form of the energy functional is in some sense better behaved when compared to the integrated square residual, due to the fact that it yields a convex optimization problem. One could also view eq. (1.2) as a shifted Tikhonov regularization or equally as a smoothed total variation norm, which is common in imaging applications [rudin1992nonlinear]. Variational forms are also sometimes used as a starting point to derive loss functions for PINNs [bai2023physics, karumuri2020simulator] and deep Ritz [yu2018deep].
@@
Equations (1.2) and (1.3) are combined to construct the training loss function:
@@
where $\eta>0$ is a regularization parameter chosen to balance contributions from the data and physics. Later we will see that the regularization parameter can be interpreted as a measure of model-form error. The field reconstruction problem is then solved by minimizing this loss function. We will refer to eq. (1.4) as a physics-regularized inverse problem.
@@
In the more difficult case with measurement noise, the field reconstruction problem can easily become ill-posed, so a Bayesian approach is desirable. Methods which go about incorporating physics into the Bayesian field reconstruction problem typically do so under a GP framework. If the PDE is linear, as in our case, it is possible to define GP priors from the physics by careful consideration of the covariance kernel. For example, this idea can be found in [raissi2017machine, bai2024gaussian] where the covariance kernel is constructed using a numerical solver for the PDE. There is also [harkonen2023gaussian], which is restricted only to linear ODEs with constant coefficients. Nonlinear PDEs can be handled by promoting the physics to the likelihood [chen2021solving] and using a standard GP prior, e.g., the square exponential kernel. The maximum a posteriori (MAP) estimate is then taken as the solution to the PDE. A similar work can be found in [chen2024sparse]. This work again treats the PDE solution as a GP prior. The solution is identified by minimizing the reproducing kernel Hilbert space (RKHS) norm of the prior covariance constrained on the PDE residuals on a predefined grid. This of course adds additional assumptions through a regularizer which enforces smoothness and may not directly represent the underlying physics. Also, in both methods only a deterministic answer is given.
There is also a broad class of probabilistic numerical methods which derive a prior over the PDE solution by taking the forcing term of the PDE as stochastic. Under this stochastic PDE (SPDE) construction, we assume a random forcing $q=\xi$ (often a white noise process) and define a prior over the field $u$ through solution of the SPDE $\mathcal{A}u=\xi$, for some linear differential operator $\mathcal{A}$. In Bayesian numerical homogenization [owhadi2015bayesian], the prior is constructed via application of the Green’s function of $\mathcal{A}$ to $\xi$ resulting in a GP whose sample paths satisfy the PDE almost surely. Later, in [cockayne2016probabilistic], this treatment was adapted for Bayesian inverse problems, focusing on identification of unknown coefficients but conceptually applicable to source identification. A related approach is the statistical finite element method [girolami2021statistical], which treats both the coefficients and the source term as GPs. Along similar lines the widely adopted formulation in [lindgren2011explicit] uses SPDEs of the form $(\kappa^{2}-\Delta)^{-\alpha/2}u=\xi$ to represent a broad family of Whittle-Matérn Gaussian fields with sparse Gaussian Markov random fields. Also of note is the work in [albert2020gaussian]. Through application of Mercer’s theorem to construct a particular covariance kernel, GPs are defined whose samples are exact solutions to linear PDEs. This idea also inspired the physics-consistent neural networks as an alternative to PINNs [ranftl2023physics]. In our previous work [bilionis2016probabilistic], we constructed priors over PDE solutions by directly setting the associated Green’s function as the covariance kernel of the GP. We conjectured that this process produces a physically meaningful prior and demonstrated this computationally. This idea was adapted to the nonlinear setting following information field theory [alberts2023physics]. Later, Poot et al. revisited this approach and made a connection to the problem of capturing model-form error induced by finite element discretization [poot2024bayesian]. A recent review of related methods is presented in [poot2025bayesian].
This work builds upon our previous efforts in [bilionis2016probabilistic, alberts2023physics]. Rather than assuming a random forcing and building the prior from an SPDE, we start with the variational problem eq. (1.4) and identify the GP prior for which this objective yields the MAP estimate. This leads to a GP prior whose covariance is the Green’s function of the Poisson equation, i.e., the Brownian bridge. The resulting prior is therefore not a heuristic choice, as it arises naturally as the unique Gaussian prior consistent with the classical variational formulation of the Poisson equation.
This shift in view highlights a key conceptual difference between the two philosophies when selecting a prior in probabilistic numerics. The SPDE-based approaches enforce the PDE sample-wise so that each sample from the prior produces a solution to the PDE with random forcing. While this provides strong physical fidelity, the uncertainty is tied to the assumed forcing model and this also restricts flexibility in capturing model-form error. In contrast, our MAP estimate derived prior imposes the physics at the level of an estimator in RKHS-norm. The posterior mean will belong to the solution space $H^{1}_{0}(\Omega)$, while the sample paths need not satisfy the PDE. This avoids imposing any additional regularity assumptions on the estimator. One of the major benefits of this approach is our method’s ability to capture model-form error, and in this sense our prior can be viewed as softly enforcing the physics.
@@ 1.1 Contributions
  1. 1.
    We show the classical Dirichlet energy functional underlying the Poisson equation arises naturally as the MAP estimator of a GP regression scheme with the Brownian bridge as the prior.
  2. 2.
    We derive a finite-dimensional representation of the prior for use in applications, which places the discretization on $L^{2}(\Omega)$ rather than as test points in $\Omega$ as is typical in GP regression. This representation is provably convergent to the prior and the posterior when used in inference.
  3. 3.
    We prove that this MAP estimator, and hence the function which solves eq. (1.4), converges to the ground truth in the large-data limit. Convergence holds even in the presence of significant model-form error.
  4. 4.
    By tuning an additional hyperparameter of the prior, we connect the method to the problem of identifying model-form error. We show that this hyperparameter, which controls the prior variance, is sensitive to model-form error by enforcing the physics as a soft constraint. The hyperparameter also causes the variance of the approximation to $q$ to adjust in the context of inverse problems.
@@ 1.2 Outline
The paper is organized as follows. In Sec. 2, we provide the necessary background on Gaussian process regression and kernel ridge regression. We establish the connection between the variational problem of eq. (1.4) and the Brownian bridge GP in Sec. 3. We do so by showing the loss function is the related kernel method objective, from which we deduce it is the MAP estimate of the corresponding GP regression. In 1D, we also prove the result in the setting of infinite-dimensional Bayesian inverse problems. That is, we show eq. (1.4) is the MAP estimate of the posterior obtained when starting with the Brownian bridge as a Gaussian measure on $L^{2}([0,1])$. Some analysis of the method is explored in Sec. 4. Here, we state the regularity of the prior and establish convergence conditions for the MAP estimate. We also derive a finite-dimensional approximation to the prior. Finally, in Sec. 5, we connect the method to the problem of model-form error identification. We demonstrate that the posterior of the inverse problem adjusts according to error in the specified physical model.
@@ 2 Preliminaries
We provide some necessary background on GPs and RKHSs. In Appendix A, we also provide a background on the theory of Gaussian measures, which, while used somewhat, is not the main focus in this work.
Definition 2.1 (Reproducing kernel Hilbert space [kanagawa2018gaussian]).
Let $k$ be a symmetric, positive-definite function on $\Omega\times\Omega$. A Hilbert space $H_{k}$ on $\Omega$ equipped with inner product $\langle\cdot,\cdot\rangle_{H_{k}}$ is said to be a reproducing kernel Hilbert space if the following two properties hold:
  1. 1.
@@
Property (ii) of Def. 2.1 is called the reproducing property, and the kernel defining the RKHS is called the reproducing kernel. The RKHS is uniquely determined by the positive-definite kernel that defines it, and the reverse is also true. This results from the Moore-Aronszajn theorem [aronszajn1950theory], which states that every positive definite symmetric map $k$ is associated with a unique RKHS $H_{k}$ for which $k$ is the reproducing kernel. One can show that given a positive definite kernel $k$ and its RKHS, each $f\in H_{k}$ can be written as $f=\sum_{i=1}^{\infty}\alpha_{i}k(\cdot,x_{i})$ for some $(\alpha_{i})_{i=1}^{\infty}\subset\mathbb{R}$, $(x_{i})_{i=1}^{\infty}\subset\Omega$ and $\|f\|_{H_{k}}<\infty$, where $\|f\|_{H_{k}}^{2}\coloneqq\sum_{i,j=1}^{\infty}c_{i}c_{j}k(x_{i},x_{j})$ [kanagawa2018gaussian, Equation (6)]. It is therefore easy to verify that the functions in the RKHS have the same behavior of $k$, e.g., smoothness.
@@
$(C_{k}u)(x)\coloneqq\int k(x,x^{\prime})u(x^{\prime})dx^{\prime},\quad\>u\in L^{2}(\Omega).$ (2.1)
The assumptions on $k$ imply that $C_{k}$ is a self-adjoint, positive operator, and thus has spectral decomposition
$C_{k}u=\sum_{n\in\mathbb{N}}\lambda_{n}\langle u,\psi_{n}\rangle\psi_{n},$
where $(\lambda_{n},\psi_{n})_{n=1}^{\infty}$ is the eigensystem of $C_{k}$, i.e.
$C_{k}\psi_{n}=\lambda_{n}\psi_{n},$ (2.2)
@@
Theorem 2.1 (Mercer’s Theorem [steinwart2008support]).
Let $k:\Omega\times\Omega\to\mathbb{R}$ be a continuous, positive-definite kernel, and $C_{k}$ and $(\lambda_{n},\psi_{n})_{n=1}^{\infty}$ be as given in eq. (2.1) and eq. (2.2), respectively. Then,
@@
$H_{k}=\left\{u\in L^{2}(\Omega):\sum_{n\in\mathbb{N}}\frac{1}{\lambda_{n}}\langle u,\psi_{n}\rangle^{2}<\infty\right\},$
@@
for $u,v\in H_{k}$. Hence the RKHS-norm can be expressed as $\|u\|_{H_{k}}^{2}=\sum_{n=1}^{\infty}\lambda_{n}^{-1}\langle u,\psi_{n}\rangle^{2}$. This representation is useful to us later when constructing the RKHS associated with the physics-regularized inverse problem.
@@
Definition 2.2 (Gaussian process [kanagawa2018gaussian]).
@@
Theorem 2.2 (Theorem 3.1 [kanagawa2018gaussian]).
[…]
$\tilde{m}(x)$ $\coloneqq m(x)+k_{xX}(K_{XX}+\sigma^{2}I_{n})^{-1}(y-m_{X}),\qquad x\in\Omega$ (2.4)
$\tilde{k}(x,x^{\prime})$ $\coloneqq k(x,x^{\prime})-k_{xX}(K_{XX}+\sigma^{2}I_{n})^{-1}k_{Xx^{\prime}},\qquad x,x^{\prime}\in\Omega,$ (2.5)
@@
Kernel ridge regression (KRR), or regularized least squares [caponnetto2007optimal], is closely related to GP regression. Given data in eq. (2.3), the objective of KRR is to solve the following interpolation problem
$u^{*}=\operatorname*{arg\,min}_{u\in H_{k}}\frac{1}{n}\sum_{i=1}^{n}\left(u(x_{i})-y_{i}\right)^{2}+r\|u\|_{H_{k}}^{2},$ (2.6)
where $r\geq 0$ is the regularization parameter. The inclusion of the RKHS norm in the objective function serves as a regularizer which enforces the class of functions which fit the data, while simultaneously smoothing the fit. It is known that $u$ becomes smoother as $\|u\|_{H_{k}}$ gets smaller, see [williams2006gaussian, Section 6.2]. Specifying the kernel which defines the KRR objective eq. (2.6) effectively enforces a prior on the fit. As with the GP regression posterior mean function, the solution to eq. (2.6) is also unique:
Theorem 2.3 (Theorem 3.4 [kanagawa2018gaussian]).
Let $r>0$. Then the unique solution to eq. (2.6) is
$u^{*}(x)=k_{xX}(K_{XX}+nrI_{n})^{-1}y=\sum_{i=1}^{n}a_{i}k(x,x_{i}),\quad x\in\Omega,$
where $k_{xX}=k_{Xx}^{T}\coloneqq(k(x,x_{1}),\dots,k(x,x_{n}))^{T}$ and $(a_{1},\dots,a_{n})^{T}=(K_{XX}+nrI_{n})^{-1}d$. Further, if the matrix $K_{XX}$ is invertible, then the coefficients $a_{i}$ are unique.
In [kanagawa2018gaussian], the relationship between GP regression and KRR is discussed in great detail. In a certain sense, GP regression can be viewed as the Bayesian interpretation of KRR. Notably, under mild conditions, the KRR solution and GP posterior mean function are equivalent.
@@
Let $k:\Omega\times\Omega\to\mathbb{R}$ be a positive definite kernel, and eq. (2.3) be training data. If $\sigma^{2}=n\lambda$, then $\tilde{m}=u^{*}$, where $\tilde{m}$ is the GP posterior mean function and $u^{*}$ is the unique KRR solution, given by eq. (2.4) with $m=0$ and eq. (2.6), respectively.
The equivalence between the GP posterior mean and KRR solution helps to establish much of the behavior involved with GP regression in terms of the RKHS of the prior covariance kernel. For example, it is immediate from Proposition 2.1 that the GP posterior mean function lives in the RKHS of the prior, meaning that the behavior of the posterior mean is inherited from the specified prior covariance. The last important property we need is the fact that GP sample paths a.s. do not belong to the prior RKHS, which is a consequence of Driscol’s zero-one law [driscoll1973reproducing].
Proposition 2.2 (Corollary 4.10 [kanagawa2018gaussian]).
Let $k:\Omega\times\Omega\to\mathbb{R}$ be a positive definite kernel and $H_{k}$ be the corresponding RKHS. Let $u\sim\mathcal{GP}(m,k)$ where $m\in H_{k}$. If $H_{k}$ is infinite-dimensional, then $u\in H_{k}$ with probability 0.